Grade 12th passAlgebra

Simplify the expressions and find the value if x is equal to 2
x + 7 + 4 (x – 5)

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question it is given that x = 2
We have,
= x + 7 + 4x–20
= 5x + 7–20
Then, substitute the value of x in the equation
= (5 × 2) + 7–20
= 10 + 7–20
= 17–20
= -3

Thanks